OpenStudy (anonymous):

Regular Hexagon has 12in side, whats the Area?

OpenStudy (anonymous):

https://www.google.com/#q=area+of+hexagon

OpenStudy (anonymous):

just plug in 12 and you get 374.12 as the area

OpenStudy (anonymous):

But I need to know how to do it lol Leave pi in pi form please ?

OpenStudy (solomonzelman):

Split your hexagon into 6 equilateral triangles. the formula for area of equil. trian. A=1/4 * √3 s^2 where s is the side.

OpenStudy (anonymous):

Soo...

OpenStudy (anonymous):

144/4

OpenStudy (anonymous):

36\[\sqrt{3}\]

OpenStudy (anonymous):

so \[216\sqrt{3}\]